# Unraveling the Key Components of OOD Generalization via Diversification

### v2

@@ 6 Conclusion and Limitations
  • Appendix E proves Proposition 2 of Sec. 5.1, proving the existence of a large number of pairwise diverse hypotheses which do not generalize. A proof for a similar result in multi-class classification case is also provided.
@@
  • Appendix F.2, using agreement score, explains the experimental setup and results that demonstrates that D-BAT and DivDis find hypotheses that are not only diverse but aligned with the inductive bias of the used learning algorithm.
@@
  • Appendix H provides a detailed explanation of how to construct the training and unlabeled data of 5.3 where we show that by only changing the distribution of unlabeled data, we can influence the optimal choice of the architecture. It also contains an variant of Tab. 2 with ViT&ResNet pair, as well as an extension of the experiment with a smooth interpolation from one unlabeled dataset setting to the other, showing a linear transition where one architecture goes from the optimal performance to random-chance accuracy, and vice-versa.
@@ Appendix A Proof and Discussion of Proposition 1
The first term on the right-hand side is the mutual information between $h_{\mathrm{sp}}$ and $h_{2}^{DD}$. Minimizing mutual information on the unlabeled OOD data $D_{u}$ yields an hypothesis $h_{2}^{DD}$ that disagrees with $h_{\mathrm{sp}}$ on $\frac{N_{U}}{2}$ data points while agreeing on the other $\frac{N_{U}}{2}$ (where $N_{U}$ is the size of unlabeled OOD data). In a finite sample size, this is equivalent to the hypotheses being statistically independent.
@@
  • In general, for $0\leq r_{D_{u}}\leq 0.5$, the classification boundary of $h_{2}^{DD}$ rotates counterclockwise (starting at $h(x;\frac{\pi}{4})$ for $r_{D_{u}}=0$) as the spurious ration increases i.e. $h_{2}^{DD}=h(x;\beta(r_{D_{u}}))$, where $\beta(r):[0,0.5]\rightarrow[\frac{\pi}{4},\frac{\pi}{2}]$ is an increasing function of $r$. More precisely, $\beta(r)=\frac{\pi}{2}-\arctan(\frac{1-2r}{1-r})$. This is the solution that satisfy both losses, similarly to the previous cases. The decision line lies in the 2nd and 3rd quadrants, and, therefore classifies the labeled training data correctly. The angle can be easily derived to satisfy the constraint that $h_{2}^{DD}$ and $h_{\mathrm{sp}}$ agree only on half of the unlabeled OOD data. Since $\beta(r)$ is strictly increasing, it is only when $r_{D_{u}}=0.5$ the solution of DivDis-Seq coincides with the ground truth $h^{\star}=h(x;\frac{\pi}{2})=h(x;\beta(0.5))$.
@@ Appendix B Results for Training MLPs on 2D Task
Figure 5: Performance of diversification is highly dependent on unlabeled OOD data (2D example + MLP). The unlabeled OOD data points are not shown in the plots. Left: the labeled training data $D_{t}$, the ground truth function $h^{\star}$ and the spurious function $h_{\mathrm{sp}}$. Right: the second hypothesis for D-BAT and DivDis-Seq under different spurious ratio of unlabeled OOD data.
In this section, we investigate whether the influence of the spurious ratio of unlabeled OOD data shown in Proposition 1 still holds when the learning algorithm is more flexible. Specifically, we use the same 2D settings, described in Sec. 4 and Appendix A, but we train a multilayer perceptron (MLP) instead of the linear classifier. The MLP consist of 3 fully-connected layers (with width 40) and has ReLU as the activation function.
As shown in Fig. 5, we observe that D-BAT (Pagliardini et al., 2023) finds $h^{\star}$ only in inversely correlated unlabeled OOD data, while DivDis-Seq, on the contrary, finds $h^{\star}$ under balanced unlabeled OOD data, which is consistent with the Proposition 1. Indeed, for D-BAT, when $r_{D_{u}}=0.25$ or $r_{D_{u}}=0.5$, the diversification loss contradicts with the cross-entropy loss on the labeled training data, causing misclassification on the training data. On the contrary, DivDis-Seq’s boundary rotates counterclockwise, and its diversification loss causes no contradiction with the cross-entropy training loss.
@@ Appendix C Results for DivDis on 2D Binary Task
In Sec. 4.2, we demonstrate on real-world image data that one of the key factors influencing the performance of diversification methods is the distribution of the unlabeled OOD data (more specifically, the spurious ratio $r_{D_{u}}$). Here we provide more details for the experimental setup, and results (in Fig. 9) for M/F (Pagliardini et al., 2023) dataset (where MNIST and Fashion-MNIST (Xiao et al., 2017) are concatenated). We then provide the results on CelebA (Liu et al., 2015; Lee et al., 2023) as a verification on a large-scale dataset, as shown in Tab. 4.
@@
In Fig. 10-Left, we further show results for different values of $\alpha$. We observe that tuning $\alpha$ is not sufficient to compensate for the misalignment between the unlabeled OOD data and the diversification loss, and the performance for both methods has the same trend. Specifically, larger $\alpha$ gives better test accuracy in general, as shown in Fig. 10-Left. In Fig. 10-Right, we select the best $\alpha$ for each scenario (i.e., each spurious ratio of unlabeled OOD data), and observe no meaningful difference in behaviour (compared to Fig. 2 and Fig. 9). Therefore, a conclusion similar to Proposition 1 still holds: even when tuning $\alpha$ for each unlabeled OOD data setting (i.e. spurious ratio), D-BAT performs best when the unlabeled data is inversely correlated, while DivDis performs best when the unlabeled data is balanced. This suggests that a practitioner might not be able to compensate a misalignment between unlabeled data and diversification loss by tuning the hyperparameter $\alpha$.
Results on CelebA-CC dataset. In Tab. 4, we further show results on large-scale real-world dataset, namely CelebA-CC (Liu et al., 2015; Lee et al., 2023). CelebA-CC is a variant of CelebA, introduced by (Lee et al., 2023), where the training data semantic attribute is completely correlated with the spurious attribute. Here, gender is used as the spurious attribute and hair color as the target. We take D-BAT and DivDis-Seq (for fair comparison on sequential training), and show their test accuracy on different degrees of spurious ratio of unlabeled OOD data ($r_{D_{u}}=\{0.0,0.5,1.0\}$). Consistent with our previous observations, the results show that D-BAT performs the best when $r_{D_{u}}=0.0$, and DivDis-Seq performs the best when $r_{D_{u}}=0.5$.
@@ Appendix E Proof of Proposition 2
This formulation covers our two methods of interest, D-BAT (Pagliardini et al., 2023) and DivDis (Lee et al., 2023). Indeed, the maximum agreement is upper-bounded by 0.5. For DivDis, the optimal solution has maximum agreement of 0.5, as seen in Appendix A. For D-BAT, the optimal solution has the lowest agreement possible. Indeed, for $K=2$, the optimal solution has $\mathnormal{A}_{D_{\mathrm{ood}}}(h_{1},h_{2})=0$. Thus, both methods optimal solutions are covered when upper-bounding the maximum agreement by 0.5 (as long as $K\leq(2|D_{\mathrm{ood}}|-1)$).
@@
Using a similar argument from the proof of Proposition 2, (Stepanov, 2006, 2017) tells us that for any $|D_{\mathrm{ood}}|=q^{m}$ where $m\in\mathbb{N^{+}}$, we can find a code similar to Hadamard’s with minimum distance equal to $\frac{N(q-1)}{q}$ and cardinality equal to $q^{m+1}$ = $q\cdot|D_{\mathrm{ood}}|$. By removing the semantic hypothesis from the count, we obtain that Proposition 3 holds for $K=(q\cdot|D_{\mathrm{ood}}|-1)$. ∎
@@ Appendix F Agreement score and Implicit Bias of Diverse Hypotheses
Adversarial dataset splits (Fig. 12). Atanov et al. (2022) also introduce the concept of adversarial dataset splits, which is a train-test dataset partitioning such that neural networks trained on the training set fail to generalize on the test set. To do that, they induce a spurious correlation between a high-AS discovered hypothesis $h_{D}$ and the (target) semantic hypothesis $h^{\star}$ on the training data, and the opposite correlation on the test set. Specifically, they select data points as training set $D_{t}$, such that a discovered high-AS hypothesis $h_{D}$ (specifically, $\mathrm{AS}(h_{D})>\mathrm{AS}(h^{\star})$) completely spurious correlates with $h^{\star}$, i.e. $\{x\in D_{t}:\;h_{D}(x)=h^{\star}(x)\}$. The test set $D_{\mathrm{test}}$ is constructed such that the two hypotheses are inversely correlated, i.e., $\{x\in D_{test}:\;h_{D}(x)\neq h^{\star}(x)\}$. Theoretically, a NN trained on such a training set $D_{t}$ should learn the hypotheses with a higher AS, i.e., $h_{D}$, which would lead to a low accuracy when tested on $D_{\mathrm{test}}$. This was indeed shown to hold in practice, where the test accuracy drops from $0.8$ for a random split to $0.2$ for an adversarial split.
Adversarial splits, therefore, show that neural networks favor learning the task with a higher AS (the background color in the case of Fig. 12) when there are two hypotheses that can ’explain’ the training data equally well. In this work, we refer to this ’preference’ as an alignment between the neural network and $h_{D}$. This creates a controllable testbed for studying the effect of spurious correlations on NNs training, which we also adopt in our study.
@@ F.2 Diversification Finds Hypotheses Aligned with Inductive Biases
Figure 13: Illustration of the constructed CIFAR-10 data with spurious correlation. The semantic binary labelling $h^{\star}$ is defined by a 5 vs. 5 split of the original CIFAR-10 classes. Spurious hypothesis/feature $h_{D}$ (color) is predictive in the training set $D_{t}$, and non-predictive on unlabeled OOD data $D_{u}$.
@@
Measuring the agreement score of the true or semantic hypothesis $h^{\star}$ gives us an estimate of the expected AS value of an hypothesis aligned with the inductive bias of $\mathcal{A}$ (otherwise, we wouldn’t expect $\mathcal{A}$ to be able to learn $h^{\star}$). We observe that the hypotheses found by DivDis and D-BAT have agreement scores similar to that of $h^{\star}$, indicating good alignment with the inductive biases of $\mathcal{A}$. Thus, optimizing Eq.  2, using neural networks as the learning algorithm, leads to diverse hypotheses implicitly biased towards those favored by its inductive biases. According to the definition of AS, such alignment is expected from an hypothesis found through empirical risk minimization (ERM), however it is not expected from diverse hypotheses (as defined in Eq. 2), given that the additional diversification loss could destroy this alignment. This analysis sheds light on the process by which diverse hypotheses are found, and puts an emphasis on the choice of a good learning algorithm, which is crucial, as we show in the subsequent Sec. 5.2.
@@
CIFAR-10 Waterbirds Hypothesis AS Test Acc.(%) AS WG Acc.(%) Semantic ($h^{\star}$) 0.83 100.0 0.84 100.0 Random ($h_{R}$) 0.63±${}_{\pm 0.01}$ 50.0 0.55±${}_{\pm 0.02}$ 50.0 D-BAT 0.82±${}_{\pm 0.01}$ 60.1±${}_{\pm 0.2}$ 0.89±${}_{\pm 0.01}$ 25.3±${}_{\pm 3.0}$ DivDis 0.81±${}_{\pm 0.02}$ 57.3±${}_{\pm 0.5}$ 0.91±${}_{\pm 0.04}$ 28.9±${}_{\pm 4.8}$
@@
Random ($h_{R}$) 0.55±${}_{\pm 0.02}$ 50.0
D-BAT (NP) 0.89±${}_{\pm 0.01}$ 25.3±${}_{\pm 3.0}$
D-BAT (P) 0.86±${}_{\pm 0.02}$ 59.1±${}_{\pm 1.6}$
DivDis (NP) 0.91±${}_{\pm 0.04}$ 28.9±${}_{\pm 4.8}$
DivDis (P) 0.86±${}_{\pm 0.02}$ 81.3±${}_{\pm 2.2}$
@@ G.1 Experimental details
  • DivDis’ default setting is to augment the data while training. This option was disabled to ensure fair comparison to D-BAT.
  • For their results, DivDis rebuilt the Waterbirds(Sagawa et al., 2020) dataset from scratch. On the contrary, D-BAT used the one provided by the WILDS(Koh et al., 2021) library. To ensure fair comparison, both methods were run using the latter version of the dataset.
@@
DivDis parameters. For Waterbirds variants, the optimizer is SGD, the number of epochs is 100, the learning rate is 0.001, the weight decay is 0.0001. The $\alpha$ parameter (referred as $\lambda$ in DivDis) was tuned over $\{0.1,1,10\}$. For Office-Home, the optimizer is SGD, the number of epochs is 50, the learning rate is 0.001, the weight decay is 0.0001, and the $\alpha$ parameter was tuned over $\{0.1,1,10\}$. For Camelyon17, the original best performing setting from DivDis was used.
D-BAT parameters. For Waterbirds variants and Office-Home, the optimizer is SGD, the learning rate is 0.001, the weight decay is 0.0001. Given that D-BAT optimizes sequentially, the number of epochs is an important parameter to tune. We tuned over $\text{epochs}\in\{30,100\}$ and $\alpha\in\{0.0001,0.1\}$. For Camelyon17, the original D-BAT best performing setting was used.
@@
OfficeHome ERM (D-BAT $h_{1}$) 10.4±${}_{\pm 0.9}$ 5.1±${}_{\pm 5.4}$ 20.6±${}_{\pm 1.4}$ 43.9±${}_{\pm 0.6}$ 25.2±${}_{\pm 1.5}$ 61.2±${}_{\pm 0.9}$ 55.5±${}_{\pm 2.9}$ 53.6±${}_{\pm 0.7}$ 58.3±${}_{\pm 0.3}$ 75.9±${}_{\pm 0.6}$
D-BAT $h_{2}$ 9.5±${}_{\pm 0.5}$ 8.1±${}_{\pm 5.9}$ 21.3±${}_{\pm 0.9}$ 46.1±${}_{\pm 0.5}$ 27.1±${}_{\pm 1.2}$ 61.9±${}_{\pm 0.7}$ ${}_{\pm 1.6}$ 55.5±${}_{\pm 0.7}$ 58.2±${}_{\pm 0.9}$ 79.2±${}_{\pm 0.4}$
DivDis (best) 7.0±${}_{\pm 0.1}$ 9.4±${}_{\pm 0.5}$ 23.1±${}_{\pm 1.2}$ 39.6±${}_{\pm 1.2}$ 28.3±${}_{\pm 0.8}$ 47.9±${}_{\pm 0.7}$ 10.3±${}_{\pm 1.9}$ 51.4±${}_{\pm 0.8}$ 55.9±${}_{\pm 0.6}$ 70.7±${}_{\pm 1.6}$
Waterbirds-CC ERM (D-BAT $h_{1}$) 8.6±${}_{\pm 2.1}$ 13.5±${}_{\pm 6.7}$ 7.0±${}_{\pm 0.3}$ 9.2±${}_{\pm 3.3}$ 35.6±${}_{\pm 3.8}$ 11.9±${}_{\pm 0.6}$ 9.2±${}_{\pm 0.8}$ 20.9±${}_{\pm 5.1}$ 30.1±${}_{\pm 2.3}$ 33.1±${}_{\pm 2.2}$
D-BAT $h_{2}$ ${}_{\pm 5.1}$ ${}_{\pm 6.8}$ 22.6±${}_{\pm 6.6}$ 15.4±${}_{\pm 6.6}$ 36.1±${}_{\pm 5.9}$ 55.5±${}_{\pm 2.8}$ 47.6±${}_{\pm 5.0}$ 49.7±${}_{\pm 9.8}$ 67.7±${}_{\pm 3.2}$ 57.2±${}_{\pm 3.7}$
DivDis (best) 25.0±${}_{\pm 2.7}$ 7.8±${}_{\pm 3.9}$ 20.4±${}_{\pm 16.9}$ 22.5±${}_{\pm 4.0}$ 49.4±${}_{\pm 10.3}$ 38.9±${}_{\pm 7.3}$ 11.2±${}_{\pm 4.7}$ 47.1±${}_{\pm 0.8}$ 70.5±${}_{\pm 4.7}$ 22.7±${}_{\pm 2.3}$
@@
Waterbirds-CC D-BAT (ViT-B/16 IN) 57.1±${}_{\pm 3.7}$ 45.0±${}_{\pm 1.4}$ 45.5±${}_{\pm 1.7}$ 44.5±${}_{\pm 1.8}$
DivDis (MoCO-v2) 49.4±${}_{\pm 10.3}$ 51.7±${}_{\pm 6.0}$ 49.6±${}_{\pm 8.3}$ 48.4±${}_{\pm 0.9}$
Office-Home D-BAT (ViT-MAE) 61.9±${}_{\pm 0.7}$ 62.6±${}_{\pm 0.1}$ 60.8±${}_{\pm 0.4}$ 61.7±${}_{\pm 0.7}$
DivDis (Resnet50 IN) 55.9±${}_{\pm 0.6}$ 54.6±${}_{\pm 0.1}$ 53.6±${}_{\pm 0.4}$ 53.1±${}_{\pm 0.2}$
Table 9: Increasing the number of hypotheses, while using the second-best inductive bias, does not bridge the performance gap with the best inductive bias. Best hypothesis performance for DivDis and corresponding hypothesis performance for D-BAT are reported. Second-best inductive bias chosen according to Fig. 3. Results are averaged over 3 seeds.
Figure 14: Naively increasing the number of hypotheses in DivDis is detrimental to performance. $y$-axis shows the worst-group accuracy of the best hypothesis on Waterbirds-CC. We use Resnet50 as a backbone. Pretraining is supervised pretraining on ImageNet-1k. “ERM pretrained” baseline stands for standard empirical risk minimization (ERM) without the usage of diversification. Results are averaged over 3 seeds. Preliminary investigation of this behaviour suggests that the diversification loss is not optimized well when K grows large (e.g. $K>8$) due to the averaging over all pairs of hypotheses, and that another aggregation function, e.g., maximum, could work better.

### v3

@@ 6 Conclusion and Limitations
Reproducibility Statement
In order to ensure that this work is reproducible, we have taken the following steps. In Appendix A and E, we provide proofs for each theoretical result (Proposition 1 and Proposition 2). For the experiments, in Appendix D, G, and H, we provide a complete description of the datasets, used models, and hyper-parameter settings. Additionally, all results from DivDis (Lee et al., 2023) and D-BAT (Pagliardini et al., 2023) are obtained using their respective published source code, ensuring a faithful representation of their methods. Finally, we provide the anonymized source code for the experiments performed in the paper.
@@
  • Appendix E proves Proposition 2 of Sec. 5.1, proving the existence of a large number of pairwise diverse hypotheses which do not generalize. A proof for a similar result in the multi-class classification case is also provided.
@@
  • Appendix F.2, using agreement score, explains the experimental setup and results that demonstrate that D-BAT and DivDis find hypotheses that are not only diverse but aligned with the inductive bias of the used learning algorithm.
@@
  • Appendix H provides a detailed explanation of how to construct the training and unlabeled data of 5.3 where we show that by only changing the distribution of unlabeled data, we can influence the optimal choice of the architecture. It also contains a variant of Tab. 2 with ViT&ResNet pair, as well as an extension of the experiment with a smooth interpolation from one unlabeled dataset setting to the other, showing a linear transition where one architecture goes from the optimal performance to random-chance accuracy, and vice-versa.
@@ Appendix A Proof and Discussion of Proposition 1
The first term on the right-hand side is the mutual information between $h_{\mathrm{sp}}$ and $h_{2}^{DD}$. Minimizing mutual information on the unlabeled OOD data $D_{u}$ yields a hypothesis $h_{2}^{DD}$ that disagrees with $h_{\mathrm{sp}}$ on $\frac{N_{U}}{2}$ data points while agreeing on the other $\frac{N_{U}}{2}$ (where $N_{U}$ is the size of unlabeled OOD data). In a finite sample size, this is equivalent to the hypotheses being statistically independent.
@@
  • In general, for $0\leq r_{D_{u}}\leq 0.5$, the classification boundary of $h_{2}^{DD}$ rotates counterclockwise (starting at $h(x;\frac{\pi}{4})$ for $r_{D_{u}}=0$) as the spurious ration increases i.e. $h_{2}^{DD}=h(x;\beta(r_{D_{u}}))$, where $\beta(r):[0,0.5]\rightarrow[\frac{\pi}{4},\frac{\pi}{2}]$ is an increasing function of $r$. More precisely, $\beta(r)=\frac{\pi}{2}-\arctan(\frac{1-2r}{1-r})$. This is the solution that satisfies both losses, similar to the previous cases. The decision line lies in the 2nd and 3rd quadrants, and, therefore classifies the labeled training data correctly. The angle can be easily derived to satisfy the constraint that $h_{2}^{DD}$ and $h_{\mathrm{sp}}$ agree only on half of the unlabeled OOD data. Since $\beta(r)$ is strictly increasing, it is only when $r_{D_{u}}=0.5$ the solution of DivDis-Seq coincides with the ground truth $h^{\star}=h(x;\frac{\pi}{2})=h(x;\beta(0.5))$.
@@ Appendix B Results for Training MLPs on 2D Task
Figure 5: Performance of diversification is highly dependent on unlabeled OOD data (2D example + MLP). The unlabeled OOD data points are not shown in the plots. Left: the labeled training data $D_{t}$, the ground truth function $h^{\star}$ and the spurious function $h_{\mathrm{sp}}$. Right: the second hypothesis for D-BAT and DivDis-Seq under different spurious ratios of unlabeled OOD data.
In this section, we investigate whether the influence of the spurious ratio of unlabeled OOD data shown in Proposition 1 still holds when the learning algorithm is more flexible. Specifically, we use the same 2D settings, described in Sec. 4 and Appendix A, but we train a multilayer perceptron (MLP) instead of the linear classifier. The MLP consists of 3 fully connected layers (with a width of 40) and has ReLU as the activation function.
As shown in Fig. 5, we observe that D-BAT (Pagliardini et al., 2023) finds $h^{\star}$ only in inversely correlated unlabeled OOD data, while DivDis-Seq, on the contrary, finds $h^{\star}$ under balanced unlabeled OOD data, which is consistent with the Proposition 1. Indeed, for D-BAT, when $r_{D_{u}}=0.25$ or $r_{D_{u}}=0.5$, the diversification loss contradicts the cross-entropy loss on the labeled training data, causing misclassification of the training data. On the contrary, DivDis-Seq’s boundary rotates counterclockwise, and its diversification loss causes no contradiction with the cross-entropy training loss.
@@ Appendix C Results for DivDis on 2D Binary Task
In Sec. 4.2, we demonstrate on real-world image data that one of the key factors influencing the performance of diversification methods is the distribution of the unlabeled OOD data (more specifically, the spurious ratio $r_{D_{u}}$). Here we provide more details for the experimental setup, and results (in Fig. 9) for M/F (Pagliardini et al., 2023) dataset (where MNIST and Fashion-MNIST (Xiao et al., 2017) are concatenated). We then provide the results on CelebA (Liu et al., 2015; Lee et al., 2023) as verification on a large-scale dataset, as shown in Tab. 4.
@@
In Fig. 10-Left, we further show results for different values of $\alpha$. We observe that tuning $\alpha$ is not sufficient to compensate for the misalignment between the unlabeled OOD data and the diversification loss, and the performance for both methods has the same trend. Specifically, larger $\alpha$ gives better test accuracy in general, as shown in Fig. 10-Left. In Fig. 10-Right, we select the best $\alpha$ for each scenario (i.e., each spurious ratio of unlabeled OOD data), and observe no meaningful difference in behavior (compared to Fig. 2 and Fig. 9). Therefore, a conclusion similar to Proposition 1 still holds: even when tuning $\alpha$ for each unlabeled OOD data setting (i.e. spurious ratio), D-BAT performs best when the unlabeled data is inversely correlated, while DivDis performs best when the unlabeled data is balanced. This suggests that a practitioner might not be able to compensate for a misalignment between unlabeled data and diversification loss by tuning the hyperparameter $\alpha$.
Results on CelebA-CC dataset. In Tab. 4, we further show results on a large-scale real-world dataset, namely CelebA-CC (Liu et al., 2015; Lee et al., 2023). CelebA-CC is a variant of CelebA, introduced by (Lee et al., 2023), where the training data semantic attribute is completely correlated with the spurious attribute. Here, gender is used as the spurious attribute and hair color as the target. We take D-BAT and DivDis-Seq (for fair comparison on sequential training), and show their test accuracy on different degrees of spurious ratio of unlabeled OOD data ($r_{D_{u}}=\{0.0,0.5,1.0\}$). Consistent with our previous observations, the results show that D-BAT performs the best when $r_{D_{u}}=0.0$, and DivDis-Seq performs the best when $r_{D_{u}}=0.5$.
Refer to captionRefer to caption
@@ Appendix E Proof of Proposition 2
This formulation covers our two methods of interest, D-BAT (Pagliardini et al., 2023) and DivDis (Lee et al., 2023). Indeed, the maximum agreement is upper-bounded by 0.5. For DivDis, the optimal solution has a maximum agreement of 0.5, as seen in Appendix A. For D-BAT, the optimal solution has the lowest agreement possible. Indeed, for $K=2$, the optimal solution has $\mathnormal{A}_{D_{\mathrm{ood}}}(h_{1},h_{2})=0$. Thus, both methods optimal solutions are covered when upper-bounding the maximum agreement by 0.5 (as long as $K\leq(2|D_{\mathrm{ood}}|-1)$).
@@
Using a similar argument from the proof of Proposition 2, (Stepanov, 2006; 2017) tells us that for any $|D_{\mathrm{ood}}|=q^{m}$ where $m\in\mathbb{N^{+}}$, we can find a code similar to Hadamard’s with minimum distance equal to $\frac{N(q-1)}{q}$ and cardinality equal to $q^{m+1}$ = $q\cdot|D_{\mathrm{ood}}|$. By removing the semantic hypothesis from the count, we obtain that Proposition 3 holds for $K=(q\cdot|D_{\mathrm{ood}}|-1)$. ∎
@@ Appendix F Agreement score and Implicit Bias of Diverse Hypotheses
Adversarial dataset splits (Fig. 12). Atanov et al. (2022) also introduces the concept of adversarial dataset splits, which is a train-test dataset partitioning such that neural networks trained on the training set fail to generalize on the test set. To do that, they induce a spurious correlation between a high-AS discovered hypothesis $h_{D}$ and the (target) semantic hypothesis $h^{\star}$ on the training data, and the opposite correlation on the test set. Specifically, they select data points as training set $D_{t}$, such that a discovered high-AS hypothesis $h_{D}$ (specifically, $\mathrm{AS}(h_{D})>\mathrm{AS}(h^{\star})$) completely spurious correlates with $h^{\star}$, i.e. $\{x\in D_{t}:\;h_{D}(x)=h^{\star}(x)\}$. The test set $D_{\mathrm{test}}$ is constructed such that the two hypotheses are inversely correlated, i.e., $\{x\in D_{test}:\;h_{D}(x)\neq h^{\star}(x)\}$. Theoretically, a NN trained on such a training set $D_{t}$ should learn the hypotheses with a higher AS, i.e., $h_{D}$, which would lead to a low accuracy when tested on $D_{\mathrm{test}}$. This was indeed shown to hold in practice, where the test accuracy drops from $0.8$ for a random split to $0.2$ for an adversarial split.
Adversarial splits, therefore, show that neural networks favor learning the task with a higher AS (the background color in the case of Fig. 12) when there are two hypotheses that can ’explain’ the training data equally well. In this work, we refer to this ’preference’ as an alignment between the neural network and $h_{D}$. This creates a controllable testbed for studying the effect of spurious correlations on NN training, which we also adopt in our study.
@@ F.2 Diversification Finds Hypotheses Aligned with Inductive Biases
Figure 13: Illustration of the constructed CIFAR-10 data with spurious correlation. The semantic binary labeling $h^{\star}$ is defined by a 5 vs. 5 split of the original CIFAR-10 classes. Spurious hypothesis/feature $h_{D}$ (color) is predictive in the training set $D_{t}$, and non-predictive on unlabeled OOD data $D_{u}$.
@@
Measuring the agreement score of the true or semantic hypothesis $h^{\star}$ gives us an estimate of the expected AS value of a hypothesis aligned with the inductive bias of $\mathcal{A}$ (otherwise, we wouldn’t expect $\mathcal{A}$ to be able to learn $h^{\star}$). We observe that the hypotheses found by DivDis and D-BAT have agreement scores similar to that of $h^{\star}$, indicating good alignment with the inductive biases of $\mathcal{A}$. Thus, optimizing Eq.  2, using neural networks as the learning algorithm, leads to diverse hypotheses implicitly biased towards those favored by its inductive biases. According to the definition of AS, such alignment is expected from a hypothesis found through empirical risk minimization (ERM), however, it is not expected from diverse hypotheses (as defined in Eq. 2), given that the additional diversification loss could destroy this alignment. This analysis sheds light on the process by which diverse hypotheses are found, and puts an emphasis on the choice of a good learning algorithm, which is crucial, as we show in the subsequent Sec. 5.2.
@@
CIFAR-10 Waterbirds
Hypothesis AS Test Acc.(%) AS WG Acc.(%)
Semantic ($h^{\star}$) 0.83 100.0 0.84 100.0
Random ($h_{R}$) 0.63±0.01 50.0 0.55±0.02 50.0
D-BAT 0.82±0.01 60.1±0.2 0.89±0.01 25.3±3.0
DivDis 0.81±0.02 57.3±0.5 0.91±0.04 28.9±4.8
@@
Random ($h_{R}$) 0.55±0.02 50.0
D-BAT (NP) 0.89±0.01 25.3±3.0
D-BAT (P) 0.86±0.02 59.1±1.6
DivDis (NP) 0.91±0.04 28.9±4.8
DivDis (P) 0.86±0.02 81.3±2.2
@@ G.1 Experimental details
  • DivDis’ default setting is to augment the data while training. This option was disabled to ensure a fair comparison to D-BAT.
  • For their results, DivDis rebuilt the Waterbirds(Sagawa et al., 2020) dataset from scratch. On the contrary, D-BAT used the one provided by the WILDS(Koh et al., 2021) library. To ensure a fair comparison, both methods were run using the latter version of the dataset.
@@
DivDis parameters. For Waterbirds variants, the optimizer is SGD, the number of epochs is 100, the learning rate is 0.001, the weight decay is 0.0001. The $\alpha$ parameter (referred as $\lambda$ in DivDis) was tuned over $\{0.1,1,10\}$. For Office-Home, the optimizer is SGD, the number of epochs is 50, the learning rate is 0.001, the weight decay is 0.0001, and the $\alpha$ parameter was tuned over $\{0.1,1,10\}$. For Camelyon17, the original best-performing setting from DivDis was used.
D-BAT parameters. For Waterbirds variants and Office-Home, the optimizer is SGD, the learning rate is 0.001, the weight decay is 0.0001. Given that D-BAT optimizes sequentially, the number of epochs is an important parameter to tune. We tuned over $\text{epochs}\in\{30,100\}$ and $\alpha\in\{0.0001,0.1\}$. For Camelyon17, the original D-BAT best-performing setting was used.
@@
OfficeHome ERM (D-BAT $h_{1}$) 10.4±0.9 5.1±5.4 20.6±1.4 43.9±0.6 25.2±1.5 61.2±0.9 55.5±2.9 53.6±0.7 58.3±0.3 75.9±0.6
D-BAT $h_{2}$ 9.5±0.5 8.1±5.9 21.3±0.9 46.1±0.5 27.1±1.2 61.9±0.7 61.6±1.6 55.5±0.7 58.2±0.9 79.2±0.4
DivDis (best) 7.0±0.1 9.4±0.5 23.1±1.2 39.6±1.2 28.3±0.8 47.9±0.7 10.3±1.9 51.4±0.8 55.9±0.6 70.7±1.6
Waterbirds-CC ERM (D-BAT $h_{1}$) 8.6±2.1 13.5±6.7 7.0±0.3 9.2±3.3 35.6±3.8 11.9±0.6 9.2±0.8 20.9±5.1 30.1±2.3 33.1±2.2
D-BAT $h_{2}$ 15.0±5.1 8.6±6.8 22.6±6.6 15.4±6.6 36.1±5.9 55.5±2.8 47.6±5.0 49.7±9.8 67.7±3.2 57.2±3.7
DivDis (best) 25.0±2.7 7.8±3.9 20.4±16.9 22.5±4.0 49.4±10.3 38.9±7.3 11.2±4.7 47.1±0.8 70.5±4.7 22.7±2.3
@@
Waterbirds-CC D-BAT (ViT-B/16 IN) 57.1±3.7 45.0±1.4 45.5±1.7 44.5±1.8
DivDis (MoCO-v2) 49.4±10.3 51.7±6.0 49.6±8.3 48.4±0.9
Office-Home D-BAT (ViT-MAE) 61.9±0.7 62.6±0.1 60.8±0.4 61.7±0.7
DivDis (Resnet50 IN) 55.9±0.6 54.6±0.1 53.6±0.4 53.1±0.2
Table 9: Increasing the number of hypotheses, while using the second-best inductive bias, does not bridge the performance gap with the best inductive bias. Best hypothesis performance for DivDis and the corresponding hypothesis performance for D-BAT are reported. The second-best inductive bias was chosen according to Fig. 3. Results are averaged over 3 seeds.
Figure 14: Naively increasing the number of hypotheses in DivDis is detrimental to performance. $y$-axis shows the worst-group accuracy of the best hypothesis on Waterbirds-CC. We use Resnet50 as a backbone. Pretraining is supervised pretraining on ImageNet-1k. “ERM pretrained” baseline stands for standard empirical risk minimization (ERM) without the usage of diversification. Results are averaged over 3 seeds. Preliminary investigation of this behavior suggests that the diversification loss is not optimized well when K grows large (e.g. $K>8$) due to the averaging over all pairs of hypotheses, and that another aggregation function, e.g., maximum, could work better.